Solve the system using elimination.

4x - 5y = 11
6x + 7y = 31
A. (-2, 2)
B. (10, 2)
C. (4, -5)
D. (4, 1)

Answer by etutorworld(19)

Solve the system using elimination.

4x - 5y = 11
6x + 7y = 31
We eliminate one of the variables and solve for the other variable
In this we can first eliminate "y". This can be done by multiplying the first equation by 7 and the second one by 5, we get
28x - 35y = 77
30x + 35y = 155
Adding the above two equations we get
58x = 232
Dividing by 58 on both sides we get
x = 4
To find the value of y we plug x=4 in any of the given equations
4(4) - 5y = 11
16 - 5y = 11
-5y = -5
y = 1
Therefore (x,y) = (4,1) [Option (d)]